# Twisted Superfields

### v7

@@ Introduction
Supersymmetric gauge theories have many attractive properties. For example, they are (i) holomorphic, (ii) invariant to local gauge transformations in superspace, and (iii) invariant to global translations in superspace. A difficulty with supersymmetric theories, however, is that they predict that for every particle that has been observed, there is another partner particle that has not yet been seen. As experiments probe higher and higher energies, the fact that no partner particle has been found becomes more problematic. Early on, supersymmetry practitioners asked whether some of the existing observed particles could actually be supersymmetric partners with each other. The HLS theorem [1] mostly rules out this possibility with some minor exceptions (like the Higgs boson being a slepton [2]). But even those exceptions are generally not accepted for other reasons.
The model presented in this paper incorporates the first two features of supersymmetry listed above, but not the third. Since the model is not invariant to superspace translations, it is not supersymmetric. Consequently, there is no a priori guarantee that quadratic divergences cancel. That being said, many quadratic divergences are cancelled for supersymmetric gauge theories primarily due to the fact that the theories are holomorphic and invariant to superspace gauge transformations. Also, local superspace gauge transformations turn scalar bosons into fermions, and fermions do not have quadratic divergences. Divergences may cancel in this theory for similar reasons. To that point, for nonsupersymmetric gauge theories similar to the one presented in this paper, it has been shown explicitly that quadratic divergences cancel to at least the two-loop order [3, 4].
@@
An advantage to this theory not being supersymmetric is that it can allow existing observed particles to be “superpartners” with each other (in the sense that superspace gauge transformations change them into each other). For example, in this model the partners of the gauge bosons (within the real “twisted superfield”) are quarks. Also, due to the fact that the theory is holomorphic and gauge invariant, it is argued that some of the nonperturbative phenomena of supersymmetric gauge theories may apply to the present theory.
@@ 1 U(3) x U(3) Symmetries and Fields
where $B_{2k}$ are Bernoulli numbers. Both $V$ and $\Lambda$ can be expanded in terms of component fields multiplied by U(6) matrices $T^{X}$, where the index X runs over the 36 adjoint indices of U(6). As with any unitary group, the same structure functions $f^{XYZ}$ satisfy both $\left[T^{X},T^{Y}\right]=if^{XYZ}T^{Z}$ and $\left[-T^{XT},-T^{YT}\right]=-if^{XYZ}T^{ZT}$. Since products of matrices in eq (1.4) only enter by way of commutators, a conjugate representation is available by replacing each $T^{X}$ in eq (1.4) with $\left(-T^{X}\right)^{T}$. In other words, a twisted real superfield that transforms by eq (1.2) also transforms as follows:
@@
$T_{\pm}^{A}={\tfrac{1}{\sqrt{2}}}\left(T_{1}^{A}\pm T_{2}^{A}\right)$ (1.7)
@@
$\Phi=\left(\begin{array}[]{cc}{\varphi_{1}\left(y\right)+\theta^{2}f_{1}\left(y\right)}&{-\sqrt{2}i\theta\chi\left(y\right)}\\ {-\sqrt{2}i\theta\tilde{\chi}\left(y\right)}&{\varphi_{2}\left(y\right)+\theta^{2}f_{2}\left(y\right)}\end{array}\right)$ (1.13)
@@
$\tilde{Q}_{mF}\to\tilde{Q}_{mF}e^{-i\left(\Lambda^{\prime}+q_{F}\Lambda_{+}^{0}T_{+}^{0}\right)}$ (1.17)
$Q_{mF}\to e^{i\left(\Lambda^{\prime}+q_{F}\Lambda_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
$q_{1}=-2,\,\,\,\,q_{2}=q_{3}=1,$
where $m\in\left\{1,2\right\}$. As a result, the following terms in the action are supergauge invariant:
@@
×$\hskip 11.38092pt\times\Big{(}\left(1-m^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)Q_{mF}^{{\dagger}}e^{2\left(gV^{\prime}+q_{F}g_{+}V_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
$\hskip 14.22636pt+\left(1-\tilde{m}^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)\tilde{Q}_{mF}e^{-2\left(gV^{\prime}+q_{F}g_{+}V_{+}^{0}T_{+}^{0}\right)}\tilde{Q}_{mF}^{{\dagger}}\Big{)}$
$-\sum_{m}\int d^{4}xd^{2}\theta\Bigg{(}\tilde{Q}_{m1}\left(m_{m11}+\sqrt{2}\Gamma_{m11}\Phi\right)Q_{m1}$
$+\sum_{F,F^{\prime}=2,3}\tilde{Q}_{mF}\left(m_{mFF^{\prime}}+\sqrt{2}\Gamma_{mFF^{\prime}}\Phi\right)Q_{mF^{\prime}}\Bigg{)}$
@@
To maintain gauge invariance in light of the transformations in eq (1.17), the above superpotential terms do not connect flavor 1 fields with flavor 2 or 3 fields. The $m^{2}_{mF}$ and $\tilde{m}^{2}_{mF}$ terms generate additional mass terms for the fundamental and conjugate scalars.
The theory presented above is free of gauge anomalies. There is a simple reason: for every fermion in the theory, there is another fermion in a conjugate representation with opposite Abelian charges. Since the theory is a gauge theory and is free of gauge anomalies, it is assumed that the theory is renormalizable.
@@
In [4], it was shown that a Wess-Zumino-like gauge is accessible for a theory with this kind of twisted supergauge invariance. After imposition of this Wess-Zumino gauge, the residual gauge invariance is just local spacetime gauge invariance. In the following, the fermions $\tilde{\lambda}$ and $\lambda$ will be referred to as “gauginos” despite the fact that they are in the (3,3*) representation of the gauge group, rather than the adjoint representation.
Although the Abelian field $V_{+}^{0}$ has its own, independent gauge transformation, the second Abelian field $V_{-}^{0}$ (group structure $T_{-}^{0}$) does not decouple from the gauge transformations described above. That is why the same designation $g$ was used for both the $V_{-}^{0}$ and nonAbelian couplings above. Nonetheless, in the Wess-Zumino gauge, since $V_{-}^{0}$ is an Abelian field, it can accommodate different charges multiplying its coupling constant when acting on different chiral fields. These charges (labelled as $q^{\prime}_{F}$ and $q^{\prime}_{A}$) are chosen to be:
$q^{\prime}_{1}=2,\,\,\,q^{\prime}_{2}=1,\,\,\,q^{\prime}_{3}=-1,\,\,\,q^{\prime}_{A}=-6\,,$ (1.20)
where $q^{\prime}_{A}$ is the charge multiplying $V_{-}^{0}$ when interacting with chiral adjoint fields $\Phi$. The reason for these choices is to reproduce the X17 data discussed in section 5. If instead, all charges for $V_{-}^{0}$ are chosen to be $1$, the theory is not able to reproduce all of the X17 data, but almost all of the other results of the paper still hold.
The classical theory has 48 parameters that can be adjusted classically: 2 coupling constants, 8 Abelian charges, 26 masses, 11 superpotential couplings, and a Fayet-Iliopoulos term. In the unification section of the paper, it is argued that both gauge couplings may be the same at the unification scale ($g_{+}=g$). It is also assumed that the following 5 parameters are zero classically:
$m_{23F}=m_{223}=m_{23}=\tilde{m}_{23}=0$ (1.21)
It is possible that many of the remaining parameters are also zero classically but are generated quantum mechanically.
@@ 2 Dynamical Symmetry Breaking
$-d_{+}^{0}$ $={\frac{g_{+}}{\sqrt{12}}}\left(\xi_{+}+\sum_{mF}q_{F}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)\right)$ (2.2)
$-d_{-}^{0}$ $=-{\frac{g}{\sqrt{12}}}\sum_{mF}q^{\prime}_{F}\left(-1\right)^{m}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)$
@@
The Abelian charges $q_{F}$ and $q^{\prime}_{F}$ for $V_{\pm}^{0}$ (including $d_{\pm}^{0}$) were defined in eqs (1.17) and (1.20). Just as for the $d$ auxiliary fields, the equations of motion can also be used to derive expressions for the $f$ auxiliary fields in terms of scalar fields.
@@
Due to eq (1.21) along with the assumption that $m_{21}$ and $\tilde{m}_{22}$ are small, it is assumed that the $m=2$ fundamental and conjugate scalars acquire vevs. Following precedent from Supersymmetric QCD (SQCD) [14, 13], the vevs in the $m=2$ sector are assumed to take the following form:
@@
Although $m_{21}$ and $\tilde{m}_{22}$ are assumed to be small, it is assumed that the scalar masses $\tilde{m}_{21}$ and $m_{22}$ are large compared to $\xi_{+}$. To accommodate a nontrivial minimum in the presence of these masses, the following is assumed classically:
$\bar{\tilde{\phi}}_{21}=\bar{\phi}_{22}=0\hskip 14.22636pt\textrm{classically}.$ (2.5)
@@
$\left\langle\varphi_{2}\right\rangle$ $=\frac{i\bar{\varphi}_{2}}{\sqrt{2}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{0}\\ {0}&{1}&{0}\end{array}\right)\,.$
@@
Assuming $m_{211}$, $m_{222}$ and $m_{\Phi}$ are small classically, the vev of the classical scalar potential is equal to
$2\left\langle\mathcal{V}\right\rangle$ $={\tfrac{1}{12}}g_{+}^{2}\left(\xi_{+}-2\bar{\phi}_{21}^{2}-\bar{\tilde{\phi}}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}\right)^{2}$ (2.7)
$+{\tfrac{1}{6}}g^{2}\left(2\bar{\phi}_{21}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}+{\tfrac{1}{6}}g^{2}\bar{\tilde{\phi}}_{22}^{4}$
$+{\tfrac{1}{4}}g^{2}\left(\bar{\tilde{\phi}}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}+\bar{\varphi}_{2}^{2}\right)^{2}$
$+2m_{21}^{2}\bar{\phi}_{21}^{2}+2\tilde{m}_{22}^{2}\bar{\tilde{\phi}}_{22}^{2}+m_{\Phi 2}^{2}\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}$ $=\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}\,.$
The minimum of the above classical scalar potential at the unification scale is found when the parameters and scalar vevs satisfy the following equations:
$m^{2}_{\Phi 2}=$ $m^{2}_{21}$ (2.8)
$\tilde{m}_{22}^{2}=$ $\tfrac{1}{3}g^{2}\bar{\phi}^{2}_{21}+\tfrac{1}{6}g^{2}\bar{\varphi}_{2}^{2}-\tfrac{2}{3}m^{2}_{\Phi 2}$
$\tfrac{1}{3}\xi_{+}=$ $2\bar{\phi}^{2}_{21}+\tfrac{4}{3}\bar{\tilde{\phi}}_{22}^{2}+\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}=$ $-\bar{\tilde{\phi}}_{22}^{2}-\bar{\varphi}_{2}^{2}-2m^{2}_{\Phi 2}/g^{2}\,,$
where based on the unification section, it has been assumed that $g_{+}=g$.
Theoretically, the vacuum chooses the values of all of the scalar vevs, subject to the restrictions of eq (2.8). Phenomenologically, the goal is to determine vacuum choices that would result in the observed particle spectrum and symmetries. From that point of view, all vevs and parameters except those on the left side of eq (2.8) may be chosen independently. The equations then determine the parameters on the left of eq (2.8). For the present model, three of the vevs are chosen to be at the following scales:
@@
$\begin{array}[]{|r|c|l|}\hline\cr\rm{vev}&\textrm{vev scale}&\mbox{Symmetry Breaking}\\ \hline\cr g\bar{\tilde{\phi}}_{23}&\sim 10^{15}\,\rm{GeV}&\mbox{Unification: SU(3)$\times$SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)}\\ g_{2}\bar{\phi}_{21}&\sim 10^{2}\,\rm{GeV}&\mbox{Electroweak: SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)$\times$U(1)}\\ g_{Z^{\prime}}\bar{\tilde{\phi}}_{22}&\sim 10^{-2}\,\rm{GeV}&\mbox{$Z^{\prime}$ acquires mass: SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)}\\ \hline\cr\end{array}$ (2.9)
@@
where $g_{2}$ is the Weak coupling and $g_{Z^{\prime}}$ is the $Z^{\prime}$ coupling that is driven to a very small value by a heavy right-handed neutrino, as discussed in section 4. The last equation in eq (2.8) causes the following vev to also be at the unification scale:
$g\bar{\phi}_{23}\sim 10^{15}\rm{GeV}.$ (2.10)
@@
In Supersymmetric QCD (SQCD) with fewer flavors than colors, the classical vacuum does not determine actual values of vevs, but just determines differences like $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$ in eq (2.8). But the SQCD quantum vacuum causes the vevs involved in those differences to get very large. In fact, the quantum vacuum drives them to infinity, so that for fewer flavors than colors, SQCD does not have a vacuum [14, 13]. In the appendix, it is pointed out that similar forces are at work in this theory, but small counterbalancing forces are also at work. The net result is that the quantum vacuum requires the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ to become very large, but not infinite. This is justification for why these two fields have vevs at a very large unification scale.
@@
$q_{F}g_{+}A_{+\mu}^{0}T_{+}^{0}+q^{\prime}_{F}gA_{-\mu}^{0}T_{-}^{0}+gA_{2\mu}^{8}T_{2}^{8}=g_{Y}A_{\mu}^{Y}T_{F}^{Y}+g_{Y^{\prime}}A_{\mu}^{Y^{\prime}}T_{F}^{Y^{\prime}}+g_{U}A_{\mu}^{U}T_{F}^{U},$ (2.12)
appearing in the action, the coupling constants and group matrices must satisfy:
$\left(\begin{array}[]{c}{g_{Y}T_{F}^{Y}}\\ {g_{Y^{\prime}}T_{F}^{Y^{\prime}}}\\ {g_{U}T_{F}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{q_{F}g_{+}T_{+}^{0}}\\ {q^{\prime}_{F}gT_{-}^{0}}\\ {gT_{2}^{8}}\end{array}\right).$ (2.13)
This is just a generalization of a Weinberg angle rotation. A more complete generalization could involve a third angle mixing the two massless fields, but that is not needed here. The reason that the group matrices on the left have a flavor dependence is because eq (2.12) involves the flavor-dependent Abelian charges $q_{F}$ and $q^{\prime}_{F}$.
The angles $\theta_{U}$ and $\phi_{U}$ in eq (2.13) are chosen so that $T^{Y}$ and $T^{Y^{\prime}}$ (for flavor 3) have zeros in their sixth diagonal slot, so that they get no mass contribution from $\bar{\phi}_{23}$ or $\bar{\tilde{\phi}}_{23}$. Specifically, the angles are given by:

### v8

@@ Introduction
Supersymmetric gauge theories have many attractive properties. For example, they provide a natural mechanism for cancellation of quadratic divergences and a resolution of the Hierarchy Problem. They do this by being (i) holomorphic, (ii) invariant to local gauge transformations in superspace, and (iii) invariant to global translations in superspace. A difficulty with supersymmetric theories, however, is that they predict that for every particle that has been observed, there is another partner particle that has not yet been seen. As experiments probe higher and higher energies, the fact that no partner particle has been found becomes more problematic. Early on, supersymmetry practitioners asked whether some of the existing observed particles could actually be supersymmetric partners with each other. The HLS theorem [1] mostly rules out this possibility with some minor exceptions (like the Higgs boson being a slepton [2]). But even those exceptions are generally not accepted for other reasons.
The model presented in this paper incorporates the first two features of supersymmetry listed above, but not the third. Since the model is not invariant to superspace translations, it is not supersymmetric. Consequently, there is no a priori guarantee that quadratic divergences cancel. That being said, many quadratic divergences are cancelled for supersymmetric gauge theories primarily due to the fact that the theories are holomorphic and invariant to superspace gauge transformations. Also, local superspace gauge transformations turn scalar bosons into fermions, and fermions do not have quadratic divergences. Divergences may cancel in this theory for similar reasons. To that point, for nonsupersymmetric gauge theories similar to the one presented in this paper, it has been shown explicitly that quadratic divergences cancel to at least the two-loop order [3].
@@
An advantage to this theory not being supersymmetric is that it can allow existing observed particles to be “superpartners” with each other (in the sense that superspace gauge transformations change them into each other). For example, in this model the partners of the gauge bosons (within the real “twisted superfield”) are quarks. Also, due to the fact that the theory is holomorphic and supergauge invariant, it is argued that some of the nonperturbative phenomena of supersymmetric gauge theories may apply to the present theory.
@@ 1 U(3) x U(3) Symmetries and Fields
where $B_{2k}$ are Bernoulli numbers. Both $V$ and $\Lambda$ can be expanded in terms of component fields multiplied by U(6) matrices $T^{X}$, where the index X runs over the 36 adjoint indices of U(6). As with any unitary group, the same structure functions $f^{XYZ}$ satisfy both $\left[T^{X},T^{Y}\right]=if^{XYZ}T^{Z}$ and $\left[-T^{XT},-T^{YT}\right]=-if^{XYZ}T^{ZT}$. Since products of matrices in eq (1.4) only enter by way of commutators, a conjugate representation is available by replacing each $T^{X}$ in eq (1.4) with its negative transpose $\left(-T^{X}\right)^{T}$. In other words, a twisted real superfield that transforms by eq (1.2) also transforms as follows:
@@
$T_{\pm}^{A}={\tfrac{1}{\sqrt{2}}}\left(T_{2}^{A}\pm T_{1}^{A}\right)$ (1.7)
@@
$\Phi=\left(\begin{array}[]{cc}{\varphi_{1}\left(y\right)+\theta^{2}f_{1}\left(y\right)}&{\sqrt{2}\theta\chi\left(y\right)}\\ {\sqrt{2}\theta\tilde{\chi}\left(y\right)}&{\varphi_{2}\left(y\right)+\theta^{2}f_{2}\left(y\right)}\end{array}\right)$ (1.13)
@@
$\tilde{Q}_{2F}\to\tilde{Q}_{2F}e^{-i\Lambda},\hskip 14.22636ptQ_{2F}\to e^{i\Lambda}Q_{2F},\hskip 14.22636ptF=2,3$ (1.17)
$\tilde{Q}_{21}\to\tilde{Q}_{21}e^{-i(\Lambda-\Lambda_{+}^{0}T_{+}^{0})},\hskip 14.22636ptQ_{21}\to e^{i(\Lambda-\Lambda_{+}^{0}T_{+}^{0})}Q_{21}$
$\tilde{Q}_{1F}\to\tilde{Q}_{1F}e^{-i(\Lambda-\Lambda_{+}^{0}T_{+}^{0})},\hskip 14.22636ptQ_{1F}\to e^{i(\Lambda-\Lambda_{+}^{0}T_{+}^{0})}Q_{1F}\,.$
@@
×$\hskip 11.38092pt\times\Big{(}\left(1-m^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)Q_{mF}^{{\dagger}}e^{2\left(gV^{\prime}+q_{mF}g_{+}V_{+}^{0}T_{+}^{0}\right)}Q_{mF}$
$\hskip 14.22636pt+\left(1-\tilde{m}^{2}_{mF}\theta^{2}\bar{\theta}^{2}\right)\tilde{Q}_{mF}e^{-2\left(gV^{\prime}+q_{mF}g_{+}V_{+}^{0}T_{+}^{0}\right)}\tilde{Q}_{mF}^{{\dagger}}\Big{)}$
$-\sum_{[mFF^{\prime}]}\int d^{4}xd^{2}\theta\times$
×$\hskip 11.38092pt\times\left(\tilde{Q}_{mF}\left(m_{mFF^{\prime}}+\sqrt{2}\Gamma_{mFF^{\prime}}\Phi\right)Q_{mF^{\prime}}\right)$
@@
$q_{1F}=q_{21}=-2,\,\,\,\,q_{22}=q_{23}=1\,,$
where $m\in\left\{1,2\right\}$ and $[mFF^{\prime}]$ means to only sum over combinations where $\tilde{Q}_{mF}$ and $Q_{mF^{\prime}}$ have the same $q_{mF}$ charge (and corresponding gauge transformations from eq (1.17)). The $m^{2}_{mF}$ and $\tilde{m}^{2}_{mF}$ terms generate additional mass terms for the fundamental and conjugate scalars.
The theory presented above is free of gauge anomalies. There is a simple reason: for every fermion in the theory, there is another fermion in a conjugate representation with opposite Abelian charges. Since the theory is a gauge theory and is free of gauge anomalies, the theory is renormalizable.
@@
In [3], it was shown that a Wess-Zumino-like gauge is accessible for a theory with this kind of twisted supergauge invariance. After imposition of this Wess-Zumino gauge, the residual gauge invariance is just local spacetime gauge invariance. In the following, the fermions $\lambda$ and $\tilde{\lambda}$ will be referred to as “gauginos” despite the fact that they are in the $(3,\tilde{3})$ and $(\tilde{3},3)$ representations of the gauge group, rather than the adjoint representation.
Although the Abelian field $V_{+}^{0}$ has its own, independent gauge transformation, the second Abelian field $V_{-}^{0}$ (group structure $T_{-}^{0}$) does not decouple from the gauge transformations described above. That is why the same designation $g$ was used for both the $V_{-}^{0}$ and nonAbelian couplings above. Nonetheless, in the Wess-Zumino gauge, since $V_{-}^{0}$ is an Abelian field, it can accommodate different charges multiplying its coupling constant when acting on different chiral fields. This flexibility is used to make the charges of $V_{-}^{0}$ (labelled as $q^{\prime}_{mF}$)the same as those for $V_{+}^{0}$ defined in eq (1.18):
$q^{\prime}_{mF}=q_{mF}\,.$ (1.20)
The charge of $V_{-}^{0}$ is also $1$ when interacting with gauginos and the fermions in $\Phi$; it is only -2 when acting on fields with $mF=21$ or $1F$.
The classical theory has 50 parameters that can be adjusted classically: 2 coupling constants, 2 Abelian charges, 30 masses, 15 superpotential couplings, and a Fayet-Iliopoulos term. In the unification section of the paper, it is argued that both gauge couplings may be the same at the unification scale ($g_{+}=g$). It is also assumed that the following 8 parameters are zero classically:
$m_{23F}=m_{223}=\Gamma_{23F}=\Gamma_{223}=m_{23}=\tilde{m}_{23}=0$ (1.21)
Many of the remaining parameters may also be zero classically but acquire values via quantum corrections.
@@ 2 Dynamical Symmetry Breaking
$-d_{+}^{0}$ $={\frac{g_{+}}{\sqrt{12}}}\left(\xi_{+}+\sum_{mF}q_{mF}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)\right)$ (2.2)
$-d_{-}^{0}$ $=-{\frac{g}{\sqrt{12}}}\sum_{mF}q^{\prime}_{mF}\left(-1\right)^{m}\left(\phi_{mF}^{{\dagger}}\phi_{mF}-\tilde{\phi}_{mF}\tilde{\phi}_{mF}^{{\dagger}}\right)$
@@
The Abelian charges $q_{mF}$ and $q^{\prime}_{mF}$ for $V_{\pm}^{0}$ (including $d_{\pm}^{0}$) were defined in eqs (1.18) and (1.20). Just as for the $d$ auxiliary fields, the equations of motion can also be used to derive expressions for the $f$ auxiliary fields in terms of scalar fields.
@@
Due to eq (1.21) along with the assumption that $m_{21}$ and $m_{22}$ are small, it is assumed that the $m=2$ fundamental and conjugate scalars acquire vevs. Following precedent from Supersymmetric QCD (SQCD) [13, 12], the vevs in the $m=2$ sector are assumed to take the following form:
@@
Unlike $m_{21}$ and $m_{22}$, it is assumed that the scalar masses $\tilde{m}_{21}$ and $\tilde{m}_{22}$ are large compared to $\xi_{+}$. To accommodate a nontrivial minimum in the presence of these masses, the following is assumed classically:
$\bar{\tilde{\phi}}_{21}=\bar{\tilde{\phi}}_{22}=0\hskip 14.22636pt\textrm{classically}.$ (2.5)
@@
$\left\langle\varphi_{2}\right\rangle$ $=\frac{\bar{\varphi}_{2}}{\sqrt{2}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{1}\\ {0}&{0}&{0}\end{array}\right)$
[…]
@@
Assuming $m_{211}$ and $m_{222}$ are small classically, the vev of the classical scalar potential is equal to
$2\left\langle\mathcal{V}\right\rangle$ $={\tfrac{1}{12}}g_{+}^{2}\left(\xi_{+}-2\bar{\phi}_{21}^{2}+\bar{\phi}_{22}^{2}+\Delta\bar{\phi}_{3}^{2}\right)^{2}$ (2.7)
$+{\tfrac{1}{6}}g^{2}\left(2\bar{\phi}_{21}^{2}-\Delta\bar{\phi}_{3}^{2}\right)^{2}+{\tfrac{1}{6}}g^{2}\bar{\phi}_{22}^{4}$
$+\tfrac{1}{4}g^{2}\left(\bar{\phi}_{22}^{2}-\Delta\bar{\phi}_{3}^{2}+\bar{\varphi}_{2}^{2}\right)^{2}$
$+2m_{21}^{2}\bar{\phi}_{21}^{2}+2m_{22}^{2}\bar{\phi}_{22}^{2}+\tilde{m}_{\Phi}^{2}\bar{\varphi}_{2}^{2}$
$\Delta\bar{\phi}_{3}^{2}$ $=\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$
$\tilde{m}^{2}_{\Phi}$ $=m^{2}_{\Phi}+m^{2}_{\Phi 2}\,.$
It can be seen that if $\bar{\phi}_{23}^{2}=\bar{\tilde{\phi}}_{23}^{2}$ but all other vevs, masses and $\xi_{+}$ vanish, then an absolute minimum of the potential is attained: $\left\langle\mathcal{V}\right\rangle=0$. In this model, it is assumed that $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ are at a unification scale of $\sim 10^{16}$ GeV, but all other vevs, masses and $\xi_{+}$ in eq (2.7) are at the electroweak scale ($\sim 10^{2}$ GeV) or smaller. In that case, deviations from the absolute minimum value $\left\langle\mathcal{V}\right\rangle=0$ only at appear at the electroweak scale.
Well below the unification scale, the superpotential is more complicated since the couplings for different groups run differently. Nonetheless, by adjusting $\xi_{+}$, $\Delta\bar{\phi}_{3}^{2}$ and certain mass relations, classical minima can still be found where the electroweak scale is primarily defined by $\bar{\phi}_{21}$, and smaller scales are defined by $\bar{\phi}_{22}$ and $\bar{\varphi}_{2}$.
In particular, a minimum exists where three of the vevs are at the following scales:
@@
$\begin{array}[]{|r|c|l|}\hline\cr\rm{vev}&\textrm{scale}&\mbox{Symmetry Breaking}\\ \hline\cr g\bar{\phi}_{23}&\sim 10^{16}\,\rm{GeV}&\mbox{Unification: SU(3)$\times$SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)}\\ g_{2}\bar{\phi}_{21}&\sim 10^{2}\,\rm{GeV}&\mbox{Electroweak: SU(3)$\times$SU(2)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)$\times$U(1)}\\ g_{Z^{\prime}}\bar{\phi}_{22}&\sim 10^{-2}\,\rm{GeV}&\mbox{$Z^{\prime}$ acquires mass: SU(3)$\times$U(1)$\times$U(1)$\to$SU(3)$\times$U(1)}\\ \hline\cr\end{array}$ (2.8)
@@
where $g_{2}$ is the Weak coupling and $g_{Z^{\prime}}$ is the $Z^{\prime}$ coupling that is driven to a very small value by an effective anomaly, as discussed in section 4. Since $\Delta\bar{\phi}_{3}^{2}$ is at the electroweak scale, but $g\bar{\tilde{\phi}}_{23}$ is at the unification scale, the following vev must also be at the unification scale:
$g\bar{\tilde{\phi}}_{23}\sim 10^{16}\rm{GeV}.$ (2.9)
@@
In Supersymmetric QCD (SQCD), the classical vacuum does not determine actual values of vevs, but just differences like $\bar{\phi}_{23}^{2}-\bar{\tilde{\phi}}_{23}^{2}$ in eq (2.7). But for fewer flavors than colors, the SQCD quantum vacuum causes the vevs involved in those differences to get very large. In fact, the quantum vacuum drives them to infinity, so that for fewer flavors than colors, SQCD does not have a vacuum [13, 12]. In the Appendix, it is pointed out that similar forces are at work in this theory. However, they can be counterbalanced by other quantum effects or even a very small mass $m_{233}$. If $m_{233}^{2}\bar{\phi}_{23}^{2}\lesssim\xi_{+}^{2}$, such a mass still allows large vevs for $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$. But it provides a counterbalance, so that the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ do not become infinite. This is justification for why these two fields have very large unification scale vevs.
@@
$q_{mF}g_{+}A_{+\mu}^{0}T_{+}^{0}+q^{\prime}_{mF}gA_{-\mu}^{0}T_{-}^{0}+gA_{2\mu}^{8}T_{2}^{8}=g_{Y}A_{\mu}^{Y}T_{mF}^{Y}+g_{Y^{\prime}}A_{\mu}^{Y^{\prime}}T_{mF}^{Y^{\prime}}+g_{U}A_{\mu}^{U}T_{mF}^{U},$ (2.11)
appearing in the action (eqs (1.18) and (1.20)), the coupling constants and group matrices must satisfy:
$\left(\begin{array}[]{c}{g_{Y}T_{mF}^{Y}}\\ {g_{Y^{\prime}}T_{mF}^{Y^{\prime}}}\\ {g_{U}T_{mF}^{U}}\end{array}\right)=\left(\begin{array}[]{ccc}{1}&{0}&{0}\\ {0}&{\cos\phi_{U}}&{\sin\phi_{U}}\\ {0}&{-\sin\phi_{U}}&{\cos\phi_{U}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{U}}&{0}&{\sin\theta_{U}}\\ {0}&{1}&{0}\\ {-\sin\theta_{U}}&{0}&{\cos\theta_{U}}\end{array}\right)\left(\begin{array}[]{c}{q_{mF}g_{+}T_{+}^{0}}\\ {q^{\prime}_{mF}gT_{-}^{0}}\\ {gT_{2}^{8}}\end{array}\right).$ (2.12)
This is just a generalization of a Weinberg angle rotation. A more complete generalization could involve a third angle mixing the two massless fields, but that is not needed here. The reason that the group matrices on the left have an $mF$ dependence is because eq (2.11) involves the Abelian charges $q_{mF}$ and $q^{\prime}_{mF}$.
The angles $\theta_{U}$ and $\phi_{U}$ in eq (2.12) are chosen so that $T^{Y}_{23}$ and $T^{Y^{\prime}}_{23}$ have zeros in their sixth diagonal slot, so that they get no mass contribution from $\bar{\phi}_{23}$ or $\bar{\tilde{\phi}}_{23}$. Specifically, the angles are given by: